OpenStudy (anonymous):

Joel bought 8 books. Some books cost $13 each and some cost $17 each. In all, he spent $128. Which system of linear equations represents the given situation? A. 4x + 4y = 8 and 17x + 13y = 128 B. x + y = 13 and x + 17y = 128 C. x + y = 17 and 13x + y = 128 D. x + y= 8 and 13x + 17y = 128

OpenStudy (anonymous):

Okay, we know A isn't right because it doesn't make sense with the question

OpenStudy (anonymous):

I want us to take a look at D.

OpenStudy (anonymous):

Some books cost 13 dollars, so it could be 13x . Some books cost 17 dollars also. so it could be 13x + 17y = 128. some of his books are from x and some are y. Since we don't know how many books are from x and y, you would have to input x + y = 8

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

So it's D ? Because it equals out and because we don't know how many were from each it had to be x + y = 8 ?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Thank you ^^

OpenStudy (anonymous):

You're welcome o^-^o do you need anymore help?